OpenStudy (anonymous):

Find the equation of the line that is/are parallel to the line 3x-y=4 and tangent to the circle x^2+y^2=9

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

this was one of my test question that i had to do today.. and i think i did it wrong.. so i want to see how exactly this question is solved...

OpenStudy (amistre64):

what was your approach?

OpenStudy (anonymous):

well i rewrite 3x-y=4 to -y=-3x+4 to y=3x-4 so the slope =3 then i just find guess some points that satisfy the equation of the circle like (3,0) and (0,3) then used the slope with those points to write the equation because center (0,0) and (3,0) give 9 so.. i did y-3=3(x-0) y=3x+3 and y-0=3(x-3).. so y=3x+9

OpenStudy (anonymous):

that wat i did since i was running out of time

OpenStudy (amistre64):

what is the derivative of the circle equation?

OpenStudy (amistre64):

or, since the circle is centered about the origin, perp the slope y = -x/3 and place that into the circle to solve for points of tangency is one approach

OpenStudy (anonymous):

2x+2xdy/dx=0 dy/dx=-2x/2 dy/dy=-x

OpenStudy (amistre64):

2x + 2y y' = 0 y' = -x/y

OpenStudy (anonymous):

so wats the equation...none of the equation i wrote was correct?

OpenStudy (amistre64):

i like my circle approach better tho ... 

OpenStudy (anonymous):

but it says that the line 3x-y=4 is tangent to the circle

OpenStudy (amistre64):

no, it doesnt ... its says the tangent lines are parallel to that line slope of 3 is fine but derivative of the circle equation is y' = -x/y y = sqrt(9-x^2)

OpenStudy (amistre64):

\[-\frac{x}{\sqrt{9-x^2}}=3\] \[-x=3\sqrt{9-x^2}\] \[x^2=9(9-x^2)\] what is x?

OpenStudy (amistre64):

10x^2 = 81 x^2 = 81/10; x=+- 9sqrt(.1)

OpenStudy (anonymous):

81-9x^2-x^2=0 -10x^2=-81 x^2=81/10 x=+or minus 9/sq.rt 10

OpenStudy (amistre64):

then plugging that in we can determine y

OpenStudy (amistre64):

x^2 + y^2 = 9 let y=-1/3x x^2 + (-1/3x)^2 = 9 x^2 + 1/9 x^2 = 9 9x^2 + x^2 = 81 .. same results

OpenStudy (anonymous):

o woo so y=-1/3x would be the equation

OpenStudy (amistre64):

well, thats just the approach that made it simpler for me the line given has a slope of 3, a line thru the center of the circle, with a perp slope, will pass thru the circle at the points of tangency ... since the circle is centered at the origin, let y = -1/3 x and see where it hits the circle ...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks for clearing this up for me

OpenStudy (amistre64):

yep